OpenStudy (anonymous):

if you have three 1bills four 5 dollar bills and two 10 dollar bills you select a bill from random.without replacing the bill, you choose a second bill at random. what is the probilbility of a five dollar bill then a 10

OpenStudy (anonymous):

2/9

OpenStudy (kropot72):

The required probability is given by \[P($5\ then\ $10)=\frac{4}{9}\times \frac{2}{8}=\frac{1}{9}\]